### v1

# How Beginning Programmers and Code LLMs (Mis)read Each Other

Abstract.
Generative AI models, specifically large language models (LLMs), have made strides towards the long-standing goal of text-to-code generation. This progress has invited numerous studies of user interaction. However, less is known about the struggles and strategies of non-experts, for whom each step of the text-to-code problem presents challenges: describing their intent in natural language, evaluating the correctness of generated code, and editing prompts when the generated code is incorrect. This paper presents a large-scale controlled study of how 120 beginning coders across three academic institutions approach writing and editing prompts. A novel experimental design allows us to target specific steps in the text-to-code process and reveals that beginners struggle with writing and editing prompts, even for problems at their skill level and when correctness is automatically determined. Our mixed-methods evaluation provides insight into student processes and perceptions with key implications for non-expert Code LLM use within and outside of education.
Figure 1. Visualization of the multi-step process of querying an large language mode of code (Code LLM). The user starts with crafting their prompt in natural language (NL). They provide the prompt to the model, which produces code. The user then assesses the correctness of the generated code. If there are errors, they must identify how to resolve them and how to edit the prompt. This continues in an iterative fashion.
1. Introduction
Computer scientists have been working towards programming in natural language for decades (Ballard and Biermann, 1979; Sammet, 1966; Heidorn, 1974), often with the goal of making programming easier for a broader set of users. Recent advances in generative AI have brought us nearer to this goal. In programming, along with fields like digital art (Ramesh et al., 2021; Rombach et al., 2022; Radford et al., 2019), creative writing (Akoury et al., 2020; Ippolito et al., 2022; Mirowski et al., 2023), and digital music (Agostinelli et al., 2023; Liu et al., 2023a), generative AI has reduced the technical skills that users need by allowing them to prompt a model with a natural language description of their desired output.
In many fields, experts have started to use generative AI to accelerate their work, including in software engineering, where large language models of code (Code LLMs) have enhanced expert programmer productivity (Peng et al., 2023; Ziegler et al., 2022; Murali et al., 2023). However, to fulfill their potential of democratizing these fields, models must be usable without extensive technical training at each stage of creation: 1) writing prompts for the model, 2) evaluating model output for quality, and 3) iteratively refining prompts when generation is unsuccessful.
Programming presents a particularly challenging domain for non-experts. Like art, computer science has evolved an extensive technical vocabulary; since generative models are trained largely on professional code, they may not work as well if users lack this vocabulary. In visual art, music, and creative writing, a user can quickly determine whether they like the generated output even if they are not an expert (embodying the cliché “I don’t know anything about art, but I know what I like”). However, this attitude does not extend to programming. It is very challenging for a non-expert to evaluate the quality of a generated program. Even when a user knows enough to determine a generated program is incorrect, they also need to understand it well enough to know what needs to change and how to update their prompt.
In order to use a Code LLM, non-experts must grapple with a multi-step process (Figure 1). First, they must have a clear understanding of what they want the code to do. This may seem trivial, but research on requirements engineering has shown that it can be challenging (Parnas and Madey, 1995). Next, the user must clearly articulate the intended behavior of the program in natural language to the model. Once the model generates code, the user must evaluate its correctness by reading it or writing tests. If the code is not correct, they must determine what has gone wrong, and update their prompt accordingly. This requires not only understanding the generated code, but also, understanding the model’s generative process. These barriers mirror well-known challenges for non-experts with end-user programming (Ko et al., 2004) and classical AI systems (Lau, 2009).
There is a growing body of work studying how non-expert programmers use AI-assisted programming systems in naturalistic settings (Prather et al., 2023; Kazemitabaar et al., 2023). However, in open-ended tasks, it is difficult to decouple the steps of the code generation process, since they feed each other: if the user fails to identify incorrect code and moves on, their editing process can’t be observed. We present results from a carefully-controlled experiment targeting two steps in the code generation process: prompt creation (How do users describe the intended program in natural language?) and prompt modification (How do users modify their prompts when a generated program is incorrect?).
One challenge in studying how non-experts use Code LLMs is selecting tasks that make sense to them. For example, replicating Barke et al. (2023)’s insightful study of experienced programmers would not be appropriate for novices, because the tasks presuppose technical knowledge. Novices have diverse goals, backgrounds, and familiarity with mathematical and computational thinking. Our solution is to target a large population of near-novices with similar experience levels: university students who have completed a single introductory computer science course (CS1). This allows us to select tasks that are conceptually familiar to them.
Our Approach
We ask whether students who have completed CS1 can effectively prompt a Code LLM to solve tasks from their previous course. In order to isolate students’ experiences in writing and editing prompts, our experiment presents tasks as input/output pairs and tests the generated code for correctness. This provides in-depth insight into the processes they develop for describing code in natural language and iteratively refining their prompts. We pose three main research questions:
  • RQ1: Can students who have completed a CS1 course effectively prompt a Code LLM to generate code for questions from their previous courses?
  • RQ2: What is the origin of student challenges with Code LLMs? Do these differ across different groups of students?
  • RQ3: What are students’ mental models of Code LLMs and how do they effect their interactions?
We find that students struggle significantly with this task, even though we pose problems tailored to their skill level and test code correctness for them. In essence, beginning programmers and current Code LLMs tend to misread each other: the Code LLM fails to generate working code based on student descriptions and students have a hard time adapting their descriptions to the model. Our study has concerning implications for democratizing programming: if these students, who already have basic skills in code explanation and understanding, struggle with this simplified task, the full natural language-to-code task—where the user has to determine correctness themselves—must be very challenging indeed for true novices. This finding also has important implications for education. Code LLMs have sparked an intense debate over the future of computing education, including claims that traditional programming training is no longer necessary (Welsh, 2022; Manjoo, [n. d.]). By contrast, our findings highlight the continuing importance of teaching students technical communication and code understanding.
Our work differentiates itself from previous work in three key ways: scale, population, and experimental design. First, we study 120 students solving 48 different programming problems. To our knowledge, no previous work has studied user interactions with Code LLMs at this scale. Second, we focus on a near-novice population with fairly uniform levels of experience, allowing us to carefully tailor tasks to their skill level. Finally, we use an experimental paradigm that allows us to isolate the prompt writing and editing aspects of the task.
2. Related Work
Our work focuses on how programmers use LLMs to turn natural language into code. Programming with natural language is a decades old idea (Miller, 1981) and has led to several ideas about bringing programming closer to how users communicate (Myers et al., 2016). For instance, Hindle et al. (2012) imagined that future language models could be effective at turning natural language to code, a prediction that has been borne out with Code LLMs.
By exploring beginner interactions with Code LLMs, our study contributes to a growing body of work on how non-experts interact with emerging automated technologies (Van Mechelen et al., 2023), ranging from automated feedback (Feldman et al., 2018; Hsu et al., 2021; Suzuki et al., 2017) to augmented reality (Holstein et al., 2018; Radu and Schneider, 2019). We situate our study within existing work on user interactions with Code LLMs below.
Experienced programmers and LLMs
We study how beginning programmers interact with a Code LLM, the same foundational technology that powers autocomplete tools such as GitHub Copilot and others (Copilot, 2023; CodeWhisperer, 2023; TabNine, 2023). These tools are promoted as productivity-boosting technology for experienced programmers. Recent in-the-wild studies and surveys indicate that these tools are popular with expert programmers, improve their self-perception of productivity, and shift their work from writing code to understanding LLM outputs (Bird et al., 2023; Murali et al., 2023; Liang et al., 2023). In contrast, our study of beginners’ interactions with a Code LLM reveals that (1) they have mixed success with writing natural language prompts, (2) and they often struggle to understand LLM-generated code.
Vaithilingam et al. (2022) present the earliest academic study of GitHub Copilot with 24 students (undergraduate–PhD) and three tasks. Their main finding is that although participants enjoyed using it, Copilot did not help them code faster or write more correct code. We design our study for less experienced participants. For example, we developed a web interface that is much simpler than a professional IDE. The same study reports that their participants often struggled to validate LLM-generated code, and we avoid this by testing generated code for our participants automatically.
Since Copilot is a general autocomplete tool, one can use it in several ways: to produce code given code, to generate documentation from code, to turn natural language into code, and so on. Grounded Copilot (Barke et al., 2023) studies experienced programmers and reports that they prefer using it to turn natural language into code (Barke et al., 2023, Section 4.2.3). Thus our study design focuses on the natural language to code task, but with beginning programmers.
Non-experts and LLMs
Like us, several researchers have considered the impact of using Code LLM for the text-to-code task with non-experts, specifically in educational settings. Our work is larger in scale than prior work (120 students from 3 institutions and 48 problems in 8 categories), which allows us to perform statistical analyses that require large sample sizes to be reliable. Moreover, our experiment design allows us to investigate key research questions that prior work has not been able to ask, such as identifying the prompting strategies that beginners use, determining how they modify prompts that do not work, and studying several factors that affect their success.
Prather et al. (2023) study 19 students using Copilot for a final project in a CS1 course: building the game Minesweeper. They found that students struggled to use Copilot, even over the course of a week. We reach a similar conclusions with our study, with $48$ problems that are much simpler than building a working video game.
Kazemitabaar et al. (2023) develop CodingSteps, a web-based Python learning environment that allows users to query Codex. The paper compares 33 participants (10–17 years old) with access to Codex to 36 students programming independently, working on the same set of 45 programming problems over several weeks. Their findings indicate that Code LLMs may benefit student learning outcomes. However, because CodingSteps presents students with expert-written problem descriptions, their results do not shed light on whether beginners can write natural language prompts independently. They report that 32% of student prompts are verbatim copies of the expert-written problem descriptions. In contrast, our study is carefully designed to avoid this problem by showing students input/output examples instead of natural language descriptions. We also investigate the strategies that students use to understand model output and modify their prompts. Kazemitabaar et al. (2023) do not address these kinds of questions, partly because their students received feedback from instructors throughout the experiment.
Promptly (Denny et al., 2023) studies 54 students writing prompts for three CS1 problems. Our substantially larger scale (120 students and 48 problems) allows us to explore research questions beyond what they study, such as the how students change their prompting strategies, and demographic factors that influence success rates. Our paper also presents a detailed analysis of LLM output, such as the kinds of errors that appear in LLM-generated code, and the impact of non-determinism on participants’ success.
Lau and Guo (2023) interviewed 20 CS1/CS2 instructors in early 2023 about their perceptions of ChatGPT and LLM technologies. They report that instructors hold a diverse set of perspectives: some wanted to “ban it” and others felt urged to integrate these technologies into curricula to prepare students for future jobs may require using LLM technology. The students in our study echo many of the concerns and desires raised by instructors in Lau and Guo (2023).
It is also possible to use language models to assist students learning to program, without having the model write code for the student. For example, Geng et al. (2022) use language models to localize type errors in OCaml, but not to correct them. Like our study, this work isolates the interaction mode in which students use Code LLMs; however, we study prompt writing and editing, while they study error detection and explanation.
Alternatives to inline code completion
Copilot and related tools suggest inline code completions, but there are other ways to interact with AI-assisted programming tools. Vaithilingam et al. (2023) present new interfaces for Visual Studio that present code changes. Liu et al. (2023b) build a new interaction model, grounded abstraction matching (Liu et al., 2023b), which targets spreadsheets and data frames, constraining the generated code to support grounding. These ideas are exciting parallel directions for Code LLM interaction in addition to the natural language prompting approach we study here.
Code LLMs beyond text-to-code
For a beginning programmer, feedback from an expert teacher or teaching assistant can be invaluable. However, access to expert feedback is limited. There is a long line of research that tries to address this shortage by developing systems that generate actionable feedback for students (Singh et al., 2013; Rocha et al., 2023; Hollingsworth, 1960; Suzuki et al., 2017; Head et al., 2017). Phung et al. (2023) show that LLMs can help build these systems and generate higher quality feedback than prior rule-based approaches. In contrast to our human experiment, they evaluate on benchmark problems. Moreover, their system is intended to help beginners write code directly, whereas our experiment focuses on prompt writing.
Another body of work focuses on automated program repair (Goues et al., 2019), which can be used to fix trivial mistakes that frustrate beginners. Traditional automated program repair systems have required significant engineering for each programming language and problem domain. Joshi et al. (2023) show that an LLM trained to generate code can be employed to repair simple coding mistakes.
Similarly, Leinonen et al. (2023a) report that Code LLMs are better at explaining code than beginning students, and Leinonen et al. (2023b) show that an LLMs explanation of a program error can be better than default error messages. This is further evidence that LLM technology may help students learn to write code directly.
Recent additional efforts include Finnie-Ansley et al. (2022), who report that Codex is remarkably good at generating code from natural language prompts from a CS1 class and several variations of the Rainfall Problem; Dakhel et al. (2022), who compare the quality of Codex-generated code to student-written code; and Babe et al. (2023), who use student-written prompts to benchmark Code LLMs. Finally, Code LLMs have applications that go beyond natural-language-to-code, and researchers are using them as building blocks for a variety of other tasks (Schäfer et al., 2023; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2023; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2023; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023; Ferdowsi et al., 2023). The aforementioned papers present new tools, benchmarks, and studies of LLM capabilities. But, they do not study users’ abilities to prompt models, which is the focus of our work.
Using LLMs for non-programming tasks
Researchers are currently exploring a wide variety of applications for LLMs beyond computational tasks. While we do not survey the full range of such work, two recent papers are particularly relevant to our task. Zamfirescu-Pereira et al. (2023) study non-experts prompting an LLM to produce recipes. Their participants actively avoided systemic testing, which we address by automating testing. Like them, we find that participants’ mental models of LLMs are very different from how they actually work. Singh et al. (2022) compare user interactions with a multimedia writing interface with LLM-generated audio, text, and image suggestions. Our post-study interview and survey was inspired by their exploration of participant’s perceptions of AI.
3. Study Design
Our work explores whether beginning programmers can effectively prompt Code LLMs. We investigate this question through a multi-institutional (Fincher et al., 2005), lab-based study, asking 120 students who completed a CS1 course to describe 8 out of 48 possible problems presented via input/output examples.
In this section, we discuss three major aspects of our study design:
  1. (1)
    Why do we use a controlled experiment?
  2. (2)
    How do we successfully present problems to students?
  3. (3)
    How do we select problems that are appropriate for students?
We discuss the logistics of implementing the study in Section 4.
3.1. Experimental Environment: In the Lab vs. In the Classroom
Studies of student interactions with programming tools can be grouped into three main categories: studies within the context of a course during the term, post-hoc analyses of educational data, or controlled, lab-based experiments. Post-hoc analyses are not currently possible, since there is a lack of available educational Code LLM data. We discuss the decision between a course-based study and lab-based study below.
There are many benefits to real-world studies conducted in a course context, including ease of access to participants and normalized educational background (Prather et al., 2023). It is easier to study how technology directly impacts learning by using it alongside instruction (Kazemitabaar et al., 2023) or as an evaluative method (Hsu et al., 2021). At the same time, these studies cannot be as easily controlled: participation may be optional (only around 12% of students chose to participate in Denny et al. (2023)); participants may explicitly be learning through the task, making it hard to compare their responses across problems (Kazemitabaar et al., 2023); and in-depth interviews are challenging to conduct.
Lab-based studies benefit from greater uniformity in observations, which facilitates statistical analysis, and longer experimental sessions. We chose a lab-based experiment because our research questions focus on the usability of Code LLMs for beginning programmers and on their processes, rather than their educational outcomes. Specifically, the process of working with a Code LLM requires multiple, interdependent steps: (1) forming an intent, (2) crafting a prompt to describe the intent, (3) evaluating the quality of the LLM-generated code, (4) editing the prompt when the code is wrong, (5) editing the code manually, or (6) giving up and writing code manually (Figure 1). Our goal was to isolate processes (2) and (4).
Our study limits user interactions in order to isolate prompt writing and editing strategies. One key feature of our paradigm is that we automatically test the generated code. In most observational studies, programmers determine on their own whether the generated code is correct. This is itself an interesting process. However, studying this aspect of Code LLM interaction comes at the cost of studying prompt editing: if a programmer mistakenly accepts incorrect code, they will move on to the next task without editing. Prather et al. (2023) report that many of their participants mistakenly accepted incorrect code. Beginning students are particularly likely to err in this way: they may struggle to understand generated code, and their lack of confidence in their own abilities may make them trust the automated system over their own judgment (an example of automation bias (SKITKA et al., 2000; Goddard et al., 2012; Gadala, 2017; De-Arteaga et al., 2020)).
Finally, a key contribution of our work is its scale: we study 120 participants across 3 institutions and 48 programming tasks, while previous studies have had fewer participants and problems. We recruit participants from three U.S. institutions: an R1 university (Northeastern University), a small liberal arts college (Oberlin College), and a women’s college (Wellesley College). This selection increases the likelihood that our findings will generalize across institutions. Our scale allows us to explore how diverse factors, such as prior non-curricular programming experience, first-generation status, and mathematics coursework, affect participant success. These kind of statistical analyses require large sample sizes and work best with even observations of participants and problems, which are challenging to obtain in course settings.
3.2. How to Describe Problems to Students: Input/Output Examples vs. Written Descriptions
A key design decision for studies of Code LLM interactions is how to present the task. In classroom environments, students are usually given instructions for what to program via written descriptions. This makes sense, given that the student’s goal is to write code. However, natural language presentation poses critical issues for our key research questions. In our study, the goal is to write natural language descriptions of problems, not to write code. A core goal is to understand how students approach the natural-language-to-code task. If the task is presented in natural language, students may simply reuse this text rather than putting the task into their own words; our results would no longer measure beginning programmer success, but instead expert description success. Prior work shows that this is a serious concern: in Kazemitabaar et al. (2023)’s study of K-12 students, up to 49% of submissions for challenging problem categories were copied from the expert-written task description.
Even if participants do not directly copy a description, its wording could influence how participants describe the task. One challenge for beginning programmers is recalling and applying technical vocabulary; presenting them with a natural language description of the task might remind them of terminology that they would not have recalled on their own. This would endanger our goal of assessing beginning programmers’ abilities to prompt code generation models, since in many natural settings, they would not have an expert description to rely on.
We therefore rely on a popular alternative for describing program behavior: input/output examples (Figure 3). Students also could reference the function name and parameter names. Our participants had taken CS1 classes where natural language descriptions are typically accompanied by input/output examples, making this a familiar way of communicating program behavior. Several CS1 courses, and some of the assignments used in our CS1 courses, go beyond this and require students to construct their own examples or even practice test-driven development (Flatt et al., 2001; Edwards, 2004). However, our study does not require students to write their own tests.
Avoiding natural language presentation is critical in order to study how beginning programmers describe problems in their own words. However, it comes with two risks. First, the input/output paradigm may increase task difficulty, since participants must identify the key pattern on their own. Although understanding natural language descriptions of coding tasks is not always easy for beginning programmers, it is likely easier than our input/output paradigm. Second, input/output examples run the risk of underspecification (Shaw et al., 1975; Gulwani et al., 2017) – there may be more than one program that performs the correct input-output mapping. To determine that the provided tests adequately described the problem, we confirmed that our provided test sets had 100% code coverage for a correct solution and performed mutation testing (Jia and Harman, 2010). We also calculated participants’ success using only the provided test cases: if the generated code passed the provided tests, it was deemed correct, ensuring that the problem presentation aligned directly with the feedback to the user.
We feel that these potential issues pose less of a risk to our key research questions than the copy/paste or word bias risks posed by a natural language presentation. Other researchers have also used an input/output presentation paradigm in studying beginner interactions with Code LLMs (Denny et al., 2023).
3.3. Problem Selection: Previously Seen Tasks vs. New Tasks
The natural language-to-code task requires participants to describe specific programming problems. Previous work exhibits varied approaches to problem selection, from a single challenging problem in Prather et al. (2023) to three simple problems in Denny et al. (2023) to a set of 45 problems in 5 categories in Kazemitabaar et al. (2023).
Our main goal was to select problems at an appropriate level for students who had completed only CS1. Since our research questions focus on student prompting processes, not learning outcomes, we chose problems at a similar level to what participants might be able to code independently. Asking students to solve new or more complex problem types increases the likelihood that the Code LLM will generate unfamiliar or difficult to understand code, making the prompt editing process more difficult. We therefore adapted problems specifically from CS1 course materials at each institution. We made small changes to facilitate input/output testing or adjust problem difficulty.
We selected 48 problems balanced across eight conceptual categories from CS1 (Figure 2), similar to Kazemitabaar et al. (2023), but with more categories and problems. Each individual problem was assigned to 20 students; we balance the experimental lists to control for ordering effects, so that each participant solved one problem in each category, and the average difficulty of each problem list was roughly the same. To facilitate difficulty and category coverage, previous CS1 instructors were asked to provide additional problems as needed. Problems such as exp (Figure 3), for instance, require students to only recognize that numbers in a list are being squared. Other problems ask students to remember complex data structures (e.g. lists, dictionaries), but not the specific Python syntax for them. We further discuss student understanding of the problems in Section  7.2.
In order to study interactions between Code LLMs and students, it is important to select problems that cannot be trivially solved by a Code LLM without any natural language description. Very common functions (for instance, shorten_url) can be solved from a function signature alone, regardless of the accompanying description. To validate our problems, we first checked that the model could not solve problems from their function/parameter names alone and, if they could, edited the names accordingly. We also solved each problem using the Code LLM to ensure that a working natural language description existed. Finally, to address the nondeterminism of Code LLMs, we ran each validation check multiple times to obtain a stable estimate of these results (§ 6.1).
Figure 2. Study overview. (1) describes the overall student trajectory through the study. We split the post survey into two sections, divided by the semi-structured interview, to delay collecting demographic information to prevent self-bias. (2) outlines the 8 problem categories (4 timed versus 4 untimed) and the 6 problems per category. Students took individual trajectories through one problem in each category, as shown by the thin arrows. (3) showcases an example trajectory for students through the problems. Students spent, on average, 42.6 minutes (SD=10.6) completing the study, with an average of 26.6 minutes (SD=9.1) on the untimed section and 15.9 minutes (SD=3.3) on the timed section.
4. Study Logistics
The previous section (§ 3) described our multi-institutional experimental design. In this section, we discuss the logistics of participant recruitment and executing the study.
(a) An example task posed to a participant. The interface displays the function name and several input/output examples. Participants write and submit a description in the text box. During our study, 85% of students who attempted this problem wrote a successful description after a single CS1 course.
(b) We run expert tests automatically and highlight ones they fail. Students are then able to either edit their description by pressing ”Try Again” or move on to another problem.
Figure 3. The Charlie the Coding Cow interface.
4.1. Charlie Interface
We built a web application for the experiment called Charlie the Coding Cow or Charlie. Charlie presents one problem per page, displaying the function signature and several input/output examples (Figure 2(a)). Participants write natural language descriptions in a text box. When they submit a description, the Charlie server prompts Codex with the function signature and their description formatted as a docstring (Figure 4). After Codex responds, Charlie shows students the Codex-generated code and displays whether it works on the given input/output examples (Figure 2(b)).
Charlie does not permit participants to edit the generated code. If the code fails, they can retry the problem or move ahead. For retry attempts, we pre-fill the text box with their last prompt to make editing easier. Finally, after every final attempt at a problem, Charlie presents two forced-choice questions with thumbs-up / thumbs-down answers: Did Charlie generate correct code? and Would you have written this code yourself?.
Each student worked with Codex to solve 3 tutorial problems and 8 main problems. We used the Charlie character to provide distance from any AI system that students might already know. This suggested a representation that was not human and not robotic. Charlie also provides visual feedback: Charlie animates a “thinking” position while Codex generates a completion and appears in different forms when the code does or does not pass all tests.
4.2. Model Choice
When we began piloting in November 2022, the most capable Code LLM was the largest Codex model from OpenAI, code-davinci-002. Although code-davinci-002 was first released in 2021, on established Python programming benchmarks, it remains as good as gpt-3.5-turbo, which is the model presently used for GitHub Copilot’s inline completions (Zhao, 2023), the free version of ChatGPT, and several other commercial products. Specifically, gpt-3.5-turbo and code-davinci-002 score 48% and 46% respectively on the HumanEval Python programming benchmark (OpenAI, 2023; Cassano et al., 2023), the most commonly used Python benchmark for Code LLMs. Since we started our study, several other LLMs have also appeared, including non-proprietary LLMs that are better for reproducibility (§ 9). The best open models perform comparably to code-davinci-002; for instance, CodeLlama (34B) achieves 48% on HumanEval (Rozière et al., 2023a). This suggests that the model that we use is as capable at code completion as newer models used in practice.
There are larger models that are more capable, such as GPT-4, which achieves a HumanEval score of 67% (OpenAI, 2023). However, GPT-4 is significantly slower and higher latency than the alternatives, and low latency is essential for LLM code completion to be acceptable to users (Murali et al., 2023); if participants have to wait more than a few seconds for the generated code, their frustration might lead them to move on rather than re-attempting the problem.
For consistency, we used the same Codex model (code-davinci-002) throughout the study. It is important to note that Code LLMs perform best when their output is sampled; consequently, the model may produce different programs for the same prompt. We generated output using best practices for hyperparameter and sampler settings (Chen et al., 2021).
4.3. Participants
We recruited 40 participants from each institution (n=120). Eligible participants were at least 18 years old, had taken CS1 at their institution between Fall 2021 and Spring 2023, and had not completed any subsequent CS courses. We recruited participants from March-July 2023 until reaching our sample size of 120. The pilot and main study received IRB approval.
Care for Participants
Our study design sought to balance obtaining accurate data with addressing potential discomforts and power dynamics. Potential discomforts for participants included frustration regarding their inability to complete a task, which could reinforce negative perceptions of self or CS. In the tutorial, we emphasized that our goal was not to evaluate their programming skills, but the collaboration with Charlie. Students were allowed to move on from a problem at any time, resulting in a variable number of attempts per problem.
We took several steps to address potential power dynamics between students and their professors. Recruitment was done through an interest form distributed by other faculty or staff. Scheduling was performed by a researcher at another institution. Finally, research sessions were never run by a professor at the same institution as the participant.
Figure 4. A diagram of the Charlie system including the front-end interface, the back-end addition of prompts to the function declaration, sending declaration + prompt to Codex via the API, running the code completion from Codex on our pre-defined tests, and returning the results to the participant in the front-end interface.
A diagram of the Charlie system including interface, back-end processing of input and connection to Codex, reprocessing data, and feedback to the user
Figure 4. A diagram of the Charlie system including the front-end interface, the back-end addition of prompts to the function declaration, sending declaration + prompt to Codex via the API, running the code completion from Codex on our pre-defined tests, and returning the results to the participant in the front-end interface.
4.4. Study Execution
The study was conducted over Zoom with audio and video recording. Participants signed informed consent material ahead of the experiment and assented at its start. They were compensated with a \$50 gift card for the estimated 75-minute study.
Figure 2 (1) outlines the full study design. Students completed 3 tutorial problems to get familiar with the interface and see some possible Codex responses. We supplied participants with a working prompt for the first tutorial problem, then gave them a difficult problem so they could see a failure, and a final easy problem to solve independently.
The main experiment consisted of 8 problems in two blocks, the first untimed, the second timed. In the second block, students were limited to 5 minutes per problem. Participants were randomly assigned experimental lists, balanced by difficulty, using a modified Latin Square design. Four authors independently assessed each problem’s difficulty; we averaged these scores and developed six roughly equal lists (Figure 2).
After the main study, students completed a two-part survey, a semi-structured interview, and an optional debriefing session (Figure 2 (1)). The semi-structured interview was interleaved between two survey blocks to mitigate question ordering and priming biases. The first series of questions in the post-survey were adapted from previous work (Körber, 2018; Bartneck et al., 2009; Wang et al., 2021; Druga and Ko, 2021; Gorson and O’Rourke, 2020; Hart and Staveland, 1988) asking about usability, trust in automation, AI perception, growth mindset, and cognitive load. The semi-structured interview asked 8 questions covering student editing processes, what they found hard or easy, how they envisioned their interactions with Charlie, and how they imagined Charlie worked. In the optional debriefing, we explained the experiment and how Code LLMs work. The second part of the survey focused on participants’ backgrounds and demographics.
In late 2022, we ran a pilot study with 19 participants to assess the study design and usability of the interface. Pilot participants were recruited from the same three institutions as in our main study, but were students who had taken more than one CS course. This small pilot allowed us to make sure the web platform was working correctly, identify any problems with specific tasks, refine our time estimates, and assess the quality of the automatic transcriptions of the interview recordings produced by otter.ai. [^1] https://web.archive.org/web/20231205001012/https://otter.ai/ During the pilot, we identified one problem with ambiguous test cases, which we changed before the main study. Pilot participants solved an average of 5.5 out of 8 problems (an Eventual Success Rate of 68.8% using the metric described in § 5.2).
Because the average pilot participant took 53 minutes, we increased the time estimate and compensation from \$30 for 60 minutes to \$50 for 75 minutes for the main study. We also added a hidden time limit to the first block of questions in case participants spent more than 50 minutes on this portion of the study; this issue never arose in the main study.
5. Analysis
This section presents the analysis framework for § 6, § 7, and § 8. We take a mixed-methods approach to this work.
5.1. Evaluation Plan
Qualitative analysis
We collected three types of data which lend themselves to qualitative analysis: (1) information about student experience and demographics, (2) free-response questions about future use of Charlie, and (3) semi-structured interview responses. We employed both inductive and deductive open coding towards consensus. Our aim was to identify common themes present in this specific dataset, rather than to develop a theory. Two researchers with previous qualitative experience conducted the analysis. We present selected quotes from the surveys and interviews throughout. Quotations have been lightly edited to remove speech errors and any identifiable information. Each participant’s quote is accompanied by a pseudonym assigned to them during data collection.
Statistical analysis
We perform statistical testing with a significance level of $\alpha$=0.05 in order to determine whether observed differences in response measures are statistically reliable. For comparisons between two groups, we use Student’s $t$-test. For comparisons between multiple groups, we perform ANOVAs; in cases where there is no natural reference group, we use Tukey HSD tests to explore pairwise differences. We report Pearson’s $r$ for correlations between continuous variables and Kendall’s $\tau$ for correlations between continuous and ordinal variables. Where we are interested in multiple potentially interacting variables, we fit linear mixed-effects models with maximal random effects for participants and problems using the lme4 package in R (Bates et al., 2015).
5.2. Measures of Success
There are several ways to measure success when evaluating the natural language to code task. The success rate is the fraction of all attempts on which the model generates a working program. Therefore, a participant who takes several attempts to solve a problem will have a lower success rate than another who succeeds in one try. We might also ask whether a participant is ever able to solve a problem; we refer to this as the eventual success rate. This metric considers only the participant’s final attempt at each assigned problem.
Although success rates measure the correctness of the code that students saw during the experiment, LLM generation is non-deterministic. [^2] Greedy generation is significantly worse for coding tasks than non-deterministic generation (Chen et al., 2021). Therefore, studying success rates can be misleading: a participant may have just been lucky with a bad prompt or unlucky with a good prompt. For this reason, we also employ an alternative metric called pass@1, which accounts for non-deterministic generation (Chen et al., 2021). Since the debut of Codex, pass@1 has become the standard metric used to evaluate LLMs on the natural language to code task, including GPT-4 (OpenAI, 2023), Code Llama (Baptiste Rozière et al., 2023), and other models (Li et al., 2023; Fried et al., 2022; Nijkamp et al., 2022).
Given a natural language prompt, pass@1 (Chen et al., 2021) is an estimate of the probability that the LLM will generate working code in one attempt. In the LLM development literature, the accepted best practice for computing pass@1 is to query the LLM 200 times for the same prompt and test every generated program (Chen et al., 2021; Fried et al., 2022; Xu et al., 2022; Rozière et al., 2023b). Sampling 200 generations for all 2,000+ prompts is very expensive with the Codex API. Instead, we use a recently released open Code LLM called StarCoder (Li et al., 2023) that is nearly as capable as the Codex model on Python benchmarks. Pass@1 with StarCoder will be slightly lower than Codex success rates because of model differences. However, pass@1 is a more stable measure of whether a prompt will succeed than success rate. We use pass@1 for the bulk of our analyses.
5.3. Positionality
We are affiliated with the three institutions where this research was conducted; we range from undergraduate students to tenured faculty. We developed the problem lists, problem difficulty ratings, and other elements of the study design within a shared educational context. Some authors are course instructors for CS1. As described in § 4.3, significant care was taken to address power dynamics between participants and researchers. Some authors also contribute to the development and evaluation of open-source Code LLMs. Overall, the potential incentives for the research team are complex, as we approach this work as both educators and researchers. We aspire to a neutral perspective on Code LLMs, while attempting to center the student experience.
This research studies students at three selective higher education institutions in the United States. Therefore, while we are able to generalize beyond a single CS curriculum, the educational context is specific: our findings may not generalize to other settings (e.g., community colleges, K-12 education) or cultural contexts.
6. RQ1: Do Students Succeed at Prompting Code LLMs with Natural Language?
Institution Mean pass@1 Success Rate Eventual Success Rate
Oberlin 0.23 25% 61%
Wellesley 0.23 25% 57%
Northeastern 0.21 23% 54%
Overall 0.22 24% 57%
(a) Mean values of different measures of success.
(b) Success rates.
(c) Eventual success rates.
(d) Attempts.
Figure 5. Basic measures of student success at the natural language to code task. Success rate is the fraction of all attempts by a participant that succeed. Eventual success rate is the fraction of last attempts at a problem by a participant that succeed. Pass@1 resamples the LLM several times to estimate the probability of success. We present these measures by institution. Figure 4(a) presents the means. Figure 4(b) and figure 4(c) show the distribution of (eventual) success rates. Eventual success rates are higher than success rates, which is to be expected: Figure 4(d) shows that many students make several attempts at a problem before an eventual success or give up.
In this section, we present how well students do on our Code LLM prompting task and address RQ1: do students succeed at prompting Code LLMs with natural language? We explore differences between students that are linked to their ability to successfully describe problems to Code LLMs.
6.1. Basic Findings
Figure 5 presents the distribution of participants’ success rates and eventual success rates. The average participant solved 4.7 out of 8 assigned problems. The mean eventual success rate (57%) is not high, and the mean success rate (24%) is even lower, since it decreases with every failed attempt. We find no significant institutional difference for either measure of success.
Participants often submitted a large number of failing attempts (Figure 4(d)): 153 problems (aggregated across participants) required three or more attempts. In fact, one participant succeeded at a problem only after 32 attempts; another gave up after 26 attempts. These results suggest that low success rates are not due to a lack of participant effort. Participants struggled to write natural language prompts for the LLM, and often achieved success only after many repeated attempts. The challenging nature of this task is supported by comments from the students themselves (§ 7.1).
6.2. Do Participants Find the Task Challenging?
Question Scale Mean
How mentally demanding was the task? Very low-¿Very high 4
How hurried was the pace of the task? Very low-¿Very high 3.3
How successful were you? Perfect-¿Failure 3.6
How insecure, stressed, or discouraged were you? Very low-¿Very high 3.1
Table 1. Mean NASA-TLX ratings (Hart and Staveland, 1988)
In the post-survey, participants completed four items of the NASA TLX (Hart and Staveland, 1988). Overall, students found the task mentally demanding (Table 1). The questions about mental demand (Q1), time pressure (Q3), and their own performance (Q4) correlate inversely with success rate. Students whose success rates were lower generally rated the task as more demanding (Kendall’s $\tau$=-0.16; $p$=0.02); were less likely to say they were successful (Kendall’s $\tau$=-0.4; $p$¡0.0001); and reported higher levels of stress and insecurity (Kendall’s $\tau$=-0.27; $p$¡0.0001).
Self-Reported Background N Mean pass@1
International 92 0.23
Domestic 27 0.22
First-generation college student 23 0.17
Not first-generation 96 0.23
Attended private high school 38 0.22
Attended public high school 76 0.22
Raised Monolingual in English 49 0.22
Raised Monolingual Not in English 27 0.22
Raised Multilingual Including English 41 0.24
Raised Multilingual Not in English 2 0.20
Table 2. Self-reported high school, language, and family background.
6.3. Who Succeeds at the Task?
Using data from the post-survey, we analyze the relationship between pass@1 rates and previous knowledge, prior programming experience, and demographics (see Table 2 for a summary of demographics). We find only two statistically reliable differences:
  • Prior programming experience: About 1/3 of participants had no programming experience outside of CS1. The remaining participants had taken pre-college programming courses (24%), were in the next CS course (21%), or had coding experience outside of classes (29%). There is a statistically reliable difference (t-test; $p$ = 0.02) in pass@1 for students who have only coded in CS1 (0.17) versus those with additional experience (0.24).
  • First-generation college students: 19.1% of participants identified as first-generation college students. We observe a statistically reliable difference in pass@1 for first-generation participants, who struggle more with the task than others (t-test; $p$=0.04).
We examined other factors, but found no significant difference in pass@1 rates:
  • Math courses: All but one participant had taken at least one college math course and half had taken 2+ courses. Single variable calculus was the most common math course. There is no statistically reliable difference between participants who had or had not taken 2+ math courses (t-test, $p$=0.42).
  • Computing intensive majors: 42% of participants were pursuing computationally intensive majors. We observe identical pass rates for both computing and non-computing majors.
  • International students: International and U.S. domestic students had similar pass@1 rates.
  • Household language: Our participants reported growing up in households where a diverse set of languages were spoken: only English (40.8%), English and other languages (34.2%), and without English (24.2%). We were surprised to find that pass@1 did not reliably vary by childhood language. However, all participants were from selective U.S. institutions that require fluency in English, regardless of childhood language exposure.
  • Public vs private high schools: 1/3 of participants attended private schools; this had no impact on pass rates.
7. RQ2: Where do student difficulties come from?
Having shown that students find it hard to prompt a Code LLM in natural language (§ 6), we explore why. In this section, we present quantitative and qualitative results that address RQ2: when students struggle with the task, where do the struggles come from? What are the most challenging aspects of the natural language to code task?
7.1. What aspects of the task do students say are hard?
In the semi-structured interview, we asked participants to reflect on challenges and issues they encountered. Three common themes emerged: difficulties in getting Charlie to understand them; issues with the generated code; and issues stemming from students’ self-reported lack of knowledge or skill (Table 3).
Thematic Codes N
Charlie Doesn’t Understand Me 89
Issues With Generated Code 59
Student Struggles 40
No Problems Mentioned 10
Issues with Study Platform 8
Issues With Experimental Design 7
Easier To Write Code Myself 7
Table 3. Thematic codes emerging from responses to What kinds of problems or issues did you run into working with Charlie?
Charlie Doesn’t Understand Me
The most commonly raised issues related to Charlie’s understanding of prompts (n=89); we divided these into subcodes. One of the most common of these was the sentiment that Charlie failed to understand good descriptions (n=23). For instance, redCoyote commented, “It was definitely difficult to have a concept of what you wanted written in your head, and then feel like you’re articulating it well, but having it not work properly.” Similarly, aquaLadybug reports feeling helpless when a good prompt didn’t succeed: “if I was saying it […] how I thought was the best way to say it, but it still wasn’t working, I had no idea where to go from there.”
Issues with Generated Code
Another major theme was issues with the generated code. Many related to perceived bugs in the generated code or difficulty debugging (26%). Students also mentioned finding the model’s randomness frustrating (8%). khakiBee was alarmed to find that resubmitting the same prompt could generate different programs, commenting “You feel like you’ve made progress, and then because it did a different thing the next time, it’s like, what do I change? I’m trying to change what I give to the cow. And then that should change what the cow is doing. But if I’m not changing anything, why is that changing?” Some students also experienced the opposite issue: despite changing their descriptions, the model generated the same incorrect function repeatedly. purpleCarp commented, “Sometimes I changed my […] description and it just repeated the code the same. And it’s just very frustrating”. This highlights the difficulty of working with stochastic models: students expect the model output to be faithful to their descriptions.
Student Struggles
Participants also reported issues stemming from their own lack of knowledge. 10% of students reported difficulty understanding a problem, and 8% reported difficulty in understanding generated code. yellowChipmunk said, “Sometimes with the code, just given my knowledge, that’s not necessarily the way I would go about coding the code. But I think to even understand it, I would have to know what the code is trying to do, which takes more time than me just trying to reword what I said”. A handful (n=4) reported that forgetting terminology made it hard to write prompts.
7.2. Which Problems Do Students Say Are Hard?
Category Mean pass@1 Mean Eventual Success Rate Student Difficulty Ranking
Sorting* 0.09 33% 1 (Hardest)
Dictionaries* 0.16 43% 2
Nested* 0.30 68% 6
Math* 0.16 54% 4
Loops 0.13 52% 3
Lists 0.18 61% 5
Conditionals 0.33 73% 7
Strings 0.26 74% 8 (Easiest)
Table 4. Pass@1 and success rates by problem category. Each category has six problems, and an equal number of students attempted each problem. The starred (*) problems were timed. Student Difficulty Ranking is done by ordering mean Eventual Success Rate from least to greatest, as that provides as measure of what percentage of students successfully solved a given task.
Some categories of CS1 problems may be harder to solve with Code LLMs, either because the concepts are difficult or because they are difficult to describe. We examine pass@1 and eventual success rate by category as well as interview responses about which problems were challenging and easy.
We find that pass@1 and eventual success rates both vary by category (Table 4). We fit a binomial mixed-effects model to pass@1, with fixed effects of category, institution, and their interaction, and random effects of problem and participant. A statistically reliable difference in pass@1 was observed only for Sorting problems, which were the most challenging ($p$=0.045). We fit a statistical model to explore whether category difficulty varied by institution and found that participants from Oberlin had lower pass@1 in the Nested category compared to other students, but that the effect was not reliable ($p$=0.063).
Interviews provide insight into their post-task perspectives. The most commonly mentioned easiest category was Math (n=21), whereas the most common for hardest was Nested (n=19), followed by Dictionaries (n=14). These do not match the ranking in Table 4, suggesting a disconnect between student performance and perceptions of difficulty.
A common theme that emerged related to the challenge of putting understanding of the problem into English (n=44). crimsonVole said, “The ones that had huge lists of like, strings, and integers, were really hard to solve, because they were really hard to describe for me.” We differentiated this code both from students’ ability to identify patterns (n=35) and their ability to write the code without Charlie (n=8). The opposite code, Easy to Describe, applied to 36 responses from the easiest question: “I felt like time ones because they’re pretty straightforward. They’re like […] exercises that we do in my Intro CS class. And so I guess it will be easier for me to word, the description or my thinking process, like I guess that might be easier.” (yelllowWeasel).
Three codes that related to student’s lack of knowledge emerged, with 27 responses (see § 7.4 for more perspectives).
7.3. What Role Does the Model Play?
LLMs can fail in surprising ways. We now explore the kinds of model failures that participants encountered.
7.3.1. Syntax errors
Contemporary Code LLMs generally produce syntactically well-formed programs. However, 5.5% of student prompts led to Python syntax errors. We manually examined and categorized them:
  • 27 generations: Codex produces degenerate, repetitive text (Holtzman et al., 2020) or Python 2 print statements. These are model failures.
  • 81 generations: Codex could not generate a complete function within the 256 token limit ($\approx$800 characters). Our problems are simple enough to be solvable in far fewer tokens, so increasing the token limit is unlikely to help.
  • 88 generations: Codex generates incomplete code after a complete function, even with standard stop tokens.
The latter two categories arise from a trade-off in system design: the first when the interface does not request enough tokens from the Code LLM; the second when it requests so many that the model generates extraneous additional code. Although these errors are infrequent, they are hard for students to deal with. In 22.4% of these cases (n=44), students gave up after seeing the syntax error.
7.3.2. When the Model Produces Different Programs From the Same Prompt
Codex is best at coding when its output is sampled (§ 4.2), but this stochasticity can frustrate students trying to modify prompts. In 107 cases (4.2%), a student submitted a prompt several times, and in most of these cases, Codex generates a new completion. A few of these are trivially different (e.g., different variable names), but most (n=86) are different functions. Some students pointed this out in the interview – beigeHalibut noted that they “usually would run a couple times, because Charlie is not very consistent with the answers. And sometimes it works. Sometimes it wouldn’t work.”
7.3.3. When the Model Produces the Same Program Despite Changes to the Prompt
When the Code LLM produces an incorrect function, and a user edits their prompt, their intent is to have the LLM produce a different—hopefully correct—function. Frustratingly, this does not necessarily happen: sometimes the model repeatedly generates the same code despite edits to the prompt. We observe many instances where this happens (104 submissions, 11% of total): it occurs in most problems (36 of 48 problems) and is encountered by a majority of students (72 of 120 students). This often leads students to give up. In fact, out of the 340 problems where students gave up, 70 were cases where the participant edited the prompt and the LLM repeatedly generated the same code.
7.4. What Do Students Do When They Encounter Unfamiliar Python?
Figure 6. An example problem, exp, which was rated differently by Oberlin and Wellesley students, likely due to the list comprehension.
Declaration, Prompt & Completion
def exp(lst, val):
Multiplyeachnumberinthelistbythe
exponentofthegivenvalue.
return [i ** val for i in lst]
Question
Is this code you would write yourself?
Student Responses
Wellesley: Yes, Oberlin: No
Code LLMs are trained on online repositories of code and may generate code using language features that students have not seen before.
Figure 6. An example problem, exp, which was rated differently by Oberlin and Wellesley students, likely due to the list comprehension.